# PseudoNeg-MAE: Self-Supervised Point Cloud Learning using Conditional Pseudo-Negative Embeddings

### v2

@@ Abstract
We propose PseudoNeg-MAE, a novel self-supervised learning framework that enhances global feature representation of point cloud masked autoencoder by making them both discriminative and sensitive to transformations. Traditional contrastive learning methods focus on achieving invariance, discarding transformation-specific information. Recent approaches incorporate transformation sensitivity by explicitly modeling relationships between original and transformed inputs. However, they report an invariant-collapse phenomenon, where the predictor degenerates into identity mappings, resulting in latent representations that have limited variation across transformations. We propose a novel loss that explicitly penalizes invariant collapse, enabling the network to capture richer transformation cues while preserving discriminative representations. PseudoNeg-MAE uses a parametric network COPE, which learns the localized displacements caused by transformations within the latent space. However, jointly training COPE with the MAE leads to undesirable trivial solutions where COPE outputs collapse to an identity. To address this, we propose a loss that uses transformation-conditioned pseudo-negatives, to penalize such trivial invariant solutions. We validate PseudoNeg-MAE on shape classification and relative pose estimation tasks, where it achieves competitive performance on the ModelNet40 and ScanObjectNN datasets under challenging evaluation protocols and demonstrates superior accuracy in estimating relative poses compared to supervised methods.
11footnotetext: CSIRO Robotics, DATA61, CSIRO, Australia. E-mails: firstname.lastname@csiro.au22footnotetext: School of Electrical Engineering and Robotics, Queensland University of Technology (QUT), Brisbane, Australia.33footnotetext: Australian National University
Keywords: representation learning, point clouds, self-supervised learning
1 Introduction
Self-Supervised Learning (SSL) has emerged as a crucial approach in point cloud analysis, allowing models to learn robust and generalizable representations from large volumes of unlabeled 3D data. With the growing accessibility of affordable 3D sensors and scanning devices, point clouds have become a widely used data format in applications such as autonomous driving, robotics, and augmented reality. SSL methods have achieved remarkable success in various downstream tasks involving point clouds, including 3D object classification [1, 2, 3, 4], 3D scene understanding [5, 6], and 3D part segmentation [1, 2, 3, 4]. Traditional SSL methods, particularly those based on contrastive learning, focus on learning invariant representations to input perturbations or transformations. However, enforcing strict invariance to transformations may suppress critical geometric cues, such as orientation or spatial displacement, that are essential for tasks in robotics where the structure and dynamics of the physical world must be preserved.
@@ 1 Introduction
In this work, we propose a novel self-supervised learning framework, PseudoNeg-MAE, which learns transformation-sensitive representations while preserving discriminative capacity. Our key contribution is a loss function that bootstraps pseudo negatives conditioned on randomly sampled transformations, avoiding the risk of collapse to invariant solutions. These pseudo negatives serve as reference embeddings generated using the predictor, capturing the localized variations in the embedding space caused by different transformations. By incorporating them into the training objective, we encourage the model to produce representations that are sensitive to input transformations while avoiding degenerate solutions. To achive this, we introduce the COnditional Pseudo Negatives Embedding network (COPE). COPE takes transformation parameters as input and outputs a weight used to linearly project the original embeddings. This conditional mechanism enables the generation of transformation specific pseudo negatives. We use these pseudo negatives to regularize COPE through our proposed loss function, ensuring that distinct transformations yield sufficiently different embeddings in the latent space.
@@
An advantage of our method is that it formulates relative pose estimation as an optimization problem in the embedding space, using transformation-sensitive features learned via the COPE. We introduce a novel inference algorithm that iteratively aligns source and target point cloud embeddings, achieving precise alignment even with large initial misalignments. We validate the effectiveness of PseudoNeg-MAE through comprehensive experiments on shape classification and relative pose estimation tasks. On the ModelNet40 and ScanObjectNN datasets, our method achieves state-of-the-art performance, particularly under challenging protocols. For relative pose estimation tasks, our framework excels in accurately estimating relative poses between point clouds, leading to superior performance compared to supervised baselines.
2 Related Works
@@
Self-Supervised Learning for Point Cloud: Self-Supervised Learning (SSL) has become increasingly prevalent for leveraging unlabeled data, especially in the challenging domain of 3D point clouds. The inherent disorder and lack of predefined structure in point clouds necessitate diverse strategies for extracting meaningful representations. Among these, contrastive methods [6, 9, 1, 10] exploit correspondences across different views of point clouds to establish unsupervised pretraining frameworks using InfoNCE loss [11, 12]. In parallel, recent advancements have introduced transformer-based masked autoencoder approaches, such as PointBERT [13], which decodes discrete tokens from point patches, and PointMAE [3] along with PointM2AE [4], which directly predict the masked point patches rather than tokens. Extending this paradigm, Point2Vec [2] employs the principles of Data2vec [14] to facilitate feature space reconstruction in point clouds. Diverging from earlier SSL techniques that mainly induce biases towards invariant feature learning or local relationships for reconstruction, our approach focuses on learning both invariant and equivariant representations, significantly enhancing the performance across both downstream tasks.
@@
Supervised Equivariant Networks: EPN [15] and E2PN [16] introduce SE(3)-equivariant architectures designed for point cloud analysis under supervised learning. EPN builds on KPConv [17] by incorporating group convolutions over discretized SO(3) to achieve rotation equivariance, while E2PN [16] improves efficiency by using quotient representations on $S^{2}\times\mathbb{R}^{3}$ with a permutation layer to recover full rotational structure. These models show strong performance on classification and relative pose estimation tasks but rely on labeled data. Additionally, Vector Neurons [18] represent another notable class of SO(3)-equivariant networks, using vector-valued features that transform linearly under rotation. EquivReg [19] leverages Vector Neurons for correspondence-free point cloud registration, solving for relative rotations in closed-form within the equivariant feature space. Our method instead leverages the properties of equivariance as a self-supervisory signal for representation learning with a non-equivariant backbone, allowing for flexible training without labeled data while still benefiting from transformation-aware learning. We compare against these supervised baselines to highlight the advantages of our self-supervised framework.
@@
Sensitivity to Transformation in SSL methods: Contrastive learning  [12, 20, 21] has significantly advanced self-supervised representation learning; however, enforcing strict invariance to data augmentations can limit the expressive power of the learned features, especially in robotic tasks where details about specific transformations are crucial. To overcome this limitation, recent studies [22, 23, 7, 8] have explored enhancing models by incorporating sensitivity to transformations. One approach involves auxiliary tasks that require models to predict the transformations applied to the input data [24, 23, 22, 25], which helps preserve transformations-specific information within the representations and leads to less invariant feature spaces. However, these methods often lack guarantees of a consistent mapping between input transformations and changes in the latent representations [8]. Alternative approaches [26, 27], integrate loss functions that balance invariance with sensitivity to augmentations, enabling models to remain responsive to transformations while maintaining robustness. Additionally, recent methods [7, 8] have developed models that directly map transformations in the input space to transformations in the latent space, employing external predictors to modify representations accordingly. These methods [7, 8] identify the invariant-collapse problem, where predictors degenerate into identity mappings, effectively removing essential transformation cues from the learned representations. SIE [8] addresses this by replacing the linear predictor with a transformation-conditioned hypernetwork. However, without strong regularization, the predicted transformations of SIE can still converge toward the identity, leading to limited transformation sensitivity. In contrast, we introduce a novel penalty loss based on transformation-conditioned pseudo-negatives that explicitly discourages collapse into invariant representations and promotes responsiveness to transformations. Since EquiMod [7] and SIE [8] were originally developed for the image domain, we include additional experiments in the supplementary material evaluating our method against these baselines on image-based tasks. The results confirm that our proposed penalty loss effectively mitigates invariant collapse and enhances transformation sensitivity.
3 Proposed Approach - PseudoNeg-MAE
@@
In this section, we introduce PseudoNeg-MAE, a method designed to learn feature representations that are both discriminative and sensitive to transformations. Our contrastive framework is built on the Masked AutoEncoder (MAE) transformer framework, combining the strengths of both predictive (MAE) and discriminative (contrastive) model training paradigms. As shown in Fig. 1, we model the relationship between input transformations and displacements in the embedding space using linear projections. These projections are controlled by a network, COPE, which generates transformation-specific weights. We train the main network $f$ and COPE jointly. There is a risk that COPE could collapse into the identity matrix, producing invariant representations. To avoid this, we introduce a novel loss function that uses pseudo-negatives to regularize COPE and ensure that the projections are equivariant during training.
@@
Figure 1: Overview of PseudoNeg-MAE: It encodes a point cloud $\mathbf{x}_{i}$ and its transformed counterpart $\mathbf{x}_{i}^{+}$, extracting their global representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$, respectively, using a shared transformer architecture. Simultaneously, COPE network outputs the weight $\Theta_{g}$ and $\{\Theta_{g_{r}}\}$ based on the input and random transformations $g$ and $\{g_{r}\}$. $\{\Theta_{g_{r}}\}$ are used to generate pseudo-negatives ${\{\tilde{\mathbf{z}}_{i}^{g_{r}}\}}$.
Let $f$ be a parametric network that maps an input space $\mathcal{X}$ to the unit sphere. Here $f$ is a composition of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, (i.e., $f=\mathcal{E}\circ\mathcal{D}$). Let $\mathcal{G}$ be a distribution of possible transformations in the input space $\mathcal{X}$, and $g$ denotes a transformation sampled from the distribution $\mathcal{G}$. For the given input point cloud $\mathbf{x}_{i}\in\mathcal{X}$, $g$ transforms it to a positive view $\mathbf{x}_{i}^{+}=g(\mathbf{x}_{i})$. The representations $\mathbf{z}$ of $\mathbf{x}_{i},\mathbf{x}_{i}^{+}$ is defined as $\mathbf{z}_{i}=f(\mathbf{x}_{i})\text{ and }\mathbf{z}_{i}^{+}=f(\mathbf{x}_{i}^{+})$. Both $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$ reside within the same representation space $\mathcal{Z}$ of $f$. We introduce an operator $u_{g}$ that models how the transformation $g$ affects the feature representation $\mathbf{z}_{i}$.
@@
$\mathcal{L}_{\text{align}}=$ $\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{z}_{i}^{+}\right\|_{2}^{2};\mathcal{L}_{\text{unif}}=\log\mathbb{E}_{(\mathbf{z}_{i},\mathbf{z}_{k})\sim\mathcal{Z}\times\mathcal{Z}}e^{-\left\|\mathbf{z}_{i}-\mathbf{z}_{k}\right\|_{2}^{2}/\tau};\mathcal{L}_{\text{ssl}}$ $=\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{unif}}.$ (1)
@@
These approaches introduce a modified alignment term, $\mathcal{L}_{\text{align}^{*}}=\frac{1}{N}\sum_{i=1}^{N}\left\|u_{g}(\mathbf{z}_{i})-\mathbf{z}_{i}^{+}\right\|_{2}^{2}$, where $u_{g}$ defines the relationship between the feature representations $\mathbf{z}_{i}$ and their augmented counterparts $\mathbf{z}_{i}^{+}$ under the transformation $g$. When combined with the uniformity term from Eq. (1), this alignment term helps the model avoid dimensional collapse while enhancing its sensitivity to transformations. A trivial solution to $\min_{f}\mathcal{L}_{\text{align}^{*}}$ is for all transformations $\forall g\in\mathcal{G}$ to collapse into the same embedding space (i.e., $u_{g}(\mathbf{z}_{i})=\mathbf{z}_{i}$), making the embeddings invariant to transformations. To avoid this, our method, PseudoNeg-MAE, introduces an additional constraint that produces localized displacements in the embedding space, ensuring that: $u_{g_{r}}(\mathbf{z}_{i})\neq u_{g_{s}}(\mathbf{z}_{i});\quad(g_{r},g_{s})\sim\mathcal{G}\times\mathcal{G}:g_{r}\neq g_{s}.$
Sensitivity to Transformation through Pseudo Negatives: To effectively model the relationship $u_{g}$, we introduce COnditional Pseudo Negatives Embedding (COPE) network based on hypernetworks [29, 8]. Specifically, COPE takes transformation parameter $g$ as an input and outputs the weight $\Theta_{g}=\text{COPE}(g)\in\mathbb{R}^{d\times d}$. This weight is used as a linear projection on $\mathbf{z}_{i}$ which learns the localized displacements due to the transformation $g$.

### v3

@@ Abstract
We propose Point-PNG, a novel self-supervised learning framework that generates conditional pseudo‑negatives in the latent space to learn point cloud representations that are both discriminative and transformation‑sensitive. Conventional self-supervised learning methods focus on achieving invariance, discarding transformation-specific information. Recent approaches incorporate transformation sensitivity by explicitly modeling relationships between original and transformed inputs. However, they often suffer from an invariant-collapse phenomenon, where the predictor degenerates into identity mappings, resulting in latent representations with limited variation across transformations. To address this, we propose Point-PNG that explicitly penalizes invariant collapse through pseudo-negatives generation, enabling the network to capture richer transformation cues while preserving discriminative representations. To this end, we introduce a parametric network, COnditional Pseudo-Negatives Embedding (COPE), which learns localized displacements induced by transformations within the latent space. A key challenge arises when jointly training COPE with the MAE, as it tends to converge to trivial identity mappings. To overcome this, we design a loss function based on pseudo-negatives conditioned on the transformation, which penalizes such trivial invariant solutions and enforces meaningful representation learning. We validate Point-PNG on shape classification and relative pose estimation tasks, showing competitive performance on ModelNet40 and ScanObjectNN under challenging evaluation protocols, and achieving superior accuracy in relative pose estimation compared to supervised baselines.
Index Terms:
Representation learning, Point Clouds, contrastive learning, Self-Supervised Learning
[…]
=-21pt
I Introduction
Self-Supervised Learning (SSL) has emerged as a crucial approach in point cloud analysis, allowing models to learn robust and generalizable representations from large volumes of unlabeled 3D data. With the growing accessibility of affordable 3D sensors and scanning devices, point clouds have become a widely used data format in applications such as autonomous driving, robotics, and augmented reality. SSL methods have achieved remarkable success in various downstream tasks involving point clouds, including 3D object classification [1, 2, 3, 4, 5], 3D scene understanding [6, 7, 8], and 3D part segmentation [1, 2, 3, 4]. Traditional SSL methods often focus on learning invariant representations to input perturbations or transformations. However, enforcing strict invariance to transformations may suppress critical geometric cues, such as orientation or spatial displacement, that are essential for tasks in robotics and other autonomous applications where the structure and dynamics of the physical world must be preserved.
@@ I Introduction
In this work, we propose a novel self-supervised learning framework, Point-PNG, which learns transformation-sensitive representations while preserving discriminative capacity. Our key contribution is a loss function that bootstraps pseudo-negatives conditioned on randomly sampled transformations, avoiding the risk of collapse to invariant solutions. These pseudo-negatives serve as reference embeddings generated using the predictor, capturing the localized variations in the embedding space caused by different transformations. By incorporating them into the training objective, we encourage the model to produce representations that are sensitive to input transformations while avoiding degenerate solutions. To achieve this, we introduce the COnditional Pseudo Negatives Embedding network (COPE). COPE takes transformation parameters as input and outputs a weight used to linearly project the original embeddings. This conditional mechanism enables the generation of transformation specific pseudo negatives. We use these pseudo-negatives to regularize COPE through our proposed loss function, ensuring that distinct transformations yield sufficiently different embeddings in the latent space.
@@
An advantage of our method is that it formulates relative pose estimation as an optimization problem in the embedding space, using transformation-sensitive features learned via the COPE. We introduce a novel inference algorithm that iteratively aligns source and target point cloud embeddings, achieving precise alignment even with large initial misalignments.
We validate the effectiveness of Point-PNG through comprehensive experiments on shape classification and relative pose estimation tasks. On the ModelNet40 and ScanObjectNN datasets, our method achieves state-of-the-art performance, particularly under challenging protocols. For relative pose estimation tasks, our framework excels in accurately estimating relative poses between point clouds, leading to superior performance compared to supervised baselines. For completeness, we also include a head-to-head comparison on a 3D-relevant image benchmark in which inputs are 2D renderings of 3D objects undergoing controlled 3D rotations. Although rendered in 2D, this benchmark evaluates sensitivity to 3D transformations. We compare against transformation-sensitive contrastive losses [9, 10] under the same backbone and protocol, evaluating the discriminativeness and rotation sensitivity of the learned features.
II Related Works
II-A Self-Supervised Learning for Point Cloud
@@
Self-Supervised Learning (SSL) has become increasingly prevalent for leveraging unlabeled data, especially in the challenging domain of 3D point clouds. The inherent disorder and lack of predefined structure in point clouds necessitate diverse strategies for extracting meaningful representations. Among these, contrastive methods [7, 11, 1, 12] exploit correspondences across different views of point clouds to establish unsupervised pretraining frameworks using InfoNCE loss [13, 14]. In parallel, recent advancements have introduced transformer-based masked autoencoder approaches, such as PointBERT [15], which decodes discrete tokens from point patches, PointMAE [3] along with PointM2AE [4], which directly predict the masked point patches rather than tokens and PointGPT [16], which follows auto-regressive prediction of spatially ordered point patches. PCPN [17] adopts PointMAE [3] for pretraining on procedurally generated point clouds. ExpPoint-MAE [18] compares masked autoencoding and momentum-contrast [19] pretraining for point cloud transformers, and introduces a strategic unfreezing fine-tuning schedule. Concurrently, LiDAR-specific SSL methods [20, 21] adapt masked reconstruction objective to sensing characteristics. Extending this paradigm, Point2Vec [2] employs the principles of Data2vec [22] to facilitate feature space reconstruction in point clouds. Recent works extend reconstruction across modalities, either rendering 3D into 2D with differentiable renderers [23, 24, 25] or distilling multi-view image features [26] instead our work focus on single-modality representation learning. Diverging from SSL techniques that mainly induce biases towards invariant feature learning or local relationships for reconstruction, our approach focuses on learning both invariant and equivariant representations, significantly enhancing the performance across both downstream tasks.
II-B Supervised Equivariant Networks
@@
EPN [27] and E2PN [28] introduce SE(3)-equivariant architectures designed for point cloud analysis under supervised learning. EPN builds on KPConv [29] by incorporating group convolutions over discretized SO(3) to achieve rotation equivariance, while E2PN [28] improves efficiency by using quotient representations on $S^{2}\times\mathbb{R}^{3}$ with a permutation layer to recover full rotational structure. These models show strong performance on classification and relative pose estimation tasks but rely on labeled data. Additionally, Vector Neurons [30] represent another notable class of SO(3)-equivariant networks, using vector-valued features that transform linearly under rotation. EquivReg [31] leverages Vector Neurons for correspondence-free point cloud registration, solving for relative rotations in closed-form within the equivariant feature space. Our method, by contrast, leverages equivariance as a self-supervisory signal for representation learning with a non-equivariant backbone. This enables training without labeled data while still retaining the benefits of transformation-aware learning. We compare against supervised baselines to demonstrate the advantages of our self-supervised framework.
II-C Sensitivity to Transformation in SSL methods
@@
Contrastive learning  [14, 19, 32] has significantly advanced self-supervised representation learning. However, enforcing strict invariance to data augmentations can limit the expressive power of the learned features, especially in robotic tasks where details about specific transformations are crucial. To overcome this limitation, image-based studies [33, 34, 9, 10] have explored enhancing models by incorporating sensitivity to transformations. One approach involves auxiliary tasks that require models to predict the transformations applied to the input data [35, 34, 33, 36], which helps preserve transformation-specific information within the representations and leads to less invariant feature spaces. In LiDAR point clouds, PSA-SSL [37] addresses this by adding a self-supervised bounding box regression task in addition to contrastive objective to encode pose and size. However, these methods often lack guarantees of a consistent mapping between input transformations and changes in the latent representations [10]. Alternative approaches [38, 39] integrate loss functions that balance invariance with sensitivity to augmentations, enabling models to remain responsive to transformations while maintaining robustness. Additionally, recent methods [9, 10, 40] have developed models that directly map transformations in the input space to transformations in the latent space, employing external predictors to modify representations accordingly. These methods [9, 10] identify the invariant-collapse problem, where predictors degenerate into identity mappings, effectively removing essential transformation cues from the learned representations. SIE [10] addresses this by replacing the linear predictor with a transformation-conditioned hypernetwork. However, without strong regularization, the predicted transformations of SIE can still converge toward the identity, leading to limited transformation sensitivity as shown by the proof in Appendix. In contrast, we introduce a novel penalty loss based on transformation-conditioned pseudo-negatives that explicitly discourages collapse into invariant representations and promotes responsiveness to transformations. Since EquiMod [9], SEN [40], and SIE [10] were originally developed for the image domain, we include additional experiments evaluating our method against these baselines on image-based tasks. The results confirm that our proposed penalty loss effectively mitigates invariant collapse and enhances transformation sensitivity.
III Proposed Approach
@@
In this section, we introduce Point-PNG, a method designed to learn feature representations that are both discriminative and transformation-sensitive. Our framework builds on the Masked AutoEncoder (MAE) transformer, combining the strengths of predictive (MAE) and discriminative (contrastive) training paradigms. As illustrated in Fig. 1, we model the relationship between input transformations and displacements in the embedding space using linear projections, controlled by a network COPE that generates transformation-specific weights. The main network $f$ and COPE are trained jointly. However, COPE risks collapsing into the identity matrix, which would yield invariant representations. To prevent this, we introduce a novel loss function based on pseudo-negatives that regularizes COPE and enforces equivariant projections during training.
@@
Figure 1: Overview of Point-PNG: It encodes a point cloud $\scriptstyle\mathbf{x}_{i}$ and its transformed counterpart , extracting their global representations $\scriptstyle\mathbf{z}_{i}$ and $\scriptstyle\mathbf{z}_{i}^{+}$, respectively, using a shared transformer architecture. Simultaneously, COPE network outputs the weight $\scriptstyle\Theta_{g}$ and $\scriptstyle\Theta_{g_{r}}$} based on the input and random transformations $\scriptstyle{}g$ and $\scriptstyle{}g_{r}$}. $\scriptstyle\Theta_{g_{r}}$} are used to generate pseudo-negatives $\scriptstyle\tilde{\mathbf{z}}_{i}^{\,g_{r}}$}.
[…]
@@
$\begin{split}\mathcal{L}_{\text{align}}&=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{z}_{i}^{+}\right\|_{2}^{2};\\ \mathcal{L}_{\text{unif}}&=\log\mathbb{E}_{(\mathbf{z}_{i},\mathbf{z}_{k})\sim\mathcal{Z}\times\mathcal{Z}}e^{-\left\|\mathbf{z}_{i}-\mathbf{z}_{k}\right\|_{2}^{2}/\tau};\\ \mathcal{L}_{\text{ssl}}&=\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{unif}}.\end{split}$ (1)
@@
These approaches introduce a modified alignment term, $\mathcal{L}_{\text{align}^{*}}=\frac{1}{N}\sum_{i=1}^{N}\left\|u_{g}(\mathbf{z}_{i})-\mathbf{z}_{i}^{+}\right\|_{2}^{2}$, where $u_{g}$ defines the relationship between the feature representations $\mathbf{z}_{i}$ and their augmented counterparts $\mathbf{z}_{i}^{+}$ under the transformation $g$. When combined with the uniformity term from Eq. (1), this alignment term helps the model avoid dimensional collapse while enhancing its sensitivity to transformations. A trivial solution to $\min_{f}\mathcal{L}_{\text{align}^{*}}$ is for all transformations $\forall g\in\mathcal{G}$ to collapse into the same embedding space (i.e., $u_{g}(\mathbf{z}_{i})=\mathbf{z}_{i}$), making the embeddings invariant to transformations. To avoid this, our method, Point-PNG, introduces a constraint that produces localized displacements in the embedding space, ensuring that: $u_{g_{r}}(\mathbf{z}_{i})\neq u_{g_{s}}(\mathbf{z}_{i});\quad(g_{r},g_{s})\sim\mathcal{G}\times\mathcal{G}:g_{r}\neq g_{s}.$
III-A Sensitivity to Transformation through Pseudo-Negatives
To effectively model the relationship $u_{g}$, we introduce COnditional Pseudo Negatives Embedding (COPE) network, a parameter efficient row-wise weight generation module with shared parameters. Specifically, COPE takes transformation parameter $g$ as an input and outputs the weight $\Theta_{g}=\text{COPE}(g)\in\mathbb{R}^{d\times d}$. This weight is used as a linear projection on $\mathbf{z}_{i}$, which learns the localized displacements due to the transformation $g$.